A man borrowed rs. 35000 for one-year. One part of it was taken at the rate of 12% p.a. And the remaining part at the rate of 14% p.a. (simple interest). If he had to pay rs. 4460 as interest, then find the two parts which were borrowed at different rates.
Answer by mananth(16949)

Investment Part I 12.00% per annum ----x
Investment part II 14.00% per annum ----y
The sum of the investments is $35,000.00
The sum of individual interests = $4,460.00
x + y 35000 ------------------------1
12.00% x + 14.00% y = $4,460.00
Multiply by 100
12 x + 14 y = $446,000.00--------2

Multiply (1) by -12
we get

-12 x -12 y = -420000.00

Add this to (2)

0 x 2 y = $26,000.00

divide by 2

y = $13,000.00 investment at
Balance $22,000.00 investment at
CHECK
$22,000.00 @ 12.00% $2,640.00
$13,000.00 @ 14.00% $1,820.00
Total -------------------- $4,460.00